# Misspecification-robust likelihood-free inference in high dimensions

### v4

@@ Abstract
Likelihood-free inference for simulator-based statistical models has developed rapidly from its infancy to a useful tool for practitioners. However, models with more than a handful of parameters still generally remain a challenge for the Approximate Bayesian Computation (ABC) based inference. To advance the possibilities for performing likelihood-free inference in higher dimensional parameter spaces, we introduce an extension of the popular Bayesian optimisation based approach to approximate discrepancy functions in a probabilistic manner which lends itself to an efficient exploration of the parameter space. Our approach achieves computational scalability for higher dimensional parameter spaces by using separate acquisition functions and discrepancies for subsets of the parameters. The efficient additive acquisition structure is combined with exponentiated loss-likelihood to provide a misspecification-robust characterisation of posterior distributions for subsets of model parameters. The method successfully performs computationally efficient inference in a moderately sized parameter space and compares favourably to existing modularised ABC methods. We further illustrate the potential of this approach by fitting a bacterial transmission dynamics model to a real data set, which provides biologically coherent results on strain competition in a 30-dimensional parameter space.
@@ 3.2 Bayesian Optimization
Each GP model trained for each discrepancy component has an analytically accessible predictive mean and standard deviation $\mu_{j}^{*}([\theta^{*}]_{j})$ and $\sigma_{j}^{*}([\theta^{*}]_{j})$, calculated from the model components and observed data . Through the use of the Lower Confidence Bound (LCB) acquisition function $a_{LCB}()$ (Srinivas et al., 2009) with an exploration-exploitation tradeoff parameter $\beta$, this gives rise to the following acquisition function minimum for the $j$th subset:
@@
Other acquisition functions can be used in addition to the LCB, including Expected Improvement (Jones et al., 1998), and those specifically designed for likelihood-free inference (Järvenpää et al., 2019a). The parameter $\beta$ in the LCB determines the compromise between exploration and exploitation in the optimization routine. Based on our exploratory analyses of the problems considered here we have fixed parameter $\beta=0.1$. This choice may not necessarily generalize to other models and data sets, and a wider exploration will thus be needed.
@@
i.e. whichever is larger of the minimum GP predictive mean value $\mu_{j}^{*}([\theta^{*}]_{j})$ and the minimum of the observed discrepancy values $d^{obs}_{j}$ from the simulator. Because the GP function is defined over all real numbers, the predictive mean can sometimes dip below zero, making $\underset{[\theta^{*}]_{j}}{\min}(\mu_{j}^{*}([\theta^{*}]_{j}))$ hard to interpret, so in these situations, we use the minimum observed discrepancy value $\min(d^{obs}_{j})$, as this is the smallest empirically validated quantity that will lead to numerically stable solutions. It is also possible to use a non-linear transformation (e.g. logarithmic) to avoid this problem: here we did not transform the discrepancies, to clarify the analogous roles of statistical losses and discrepancies, but it will likely help in cases of numerical instability.
@@ 3.4 Summary Statistics and Partitioning
We encourage the use of any of the above methods, or in some combination, to help derive sets of summary statistics appropriate for inference. Establishing a successful Bayesian workflow generally relies upon an iterative process of predictive checks, computation validation, and model evaluation, among other steps (Gelman et al., 2020), with “black box” inference solutions often falling outside of this paradigm. For the examples in this article, we have used summaries derived from our understanding of the simulators, but the inference methodology is agnostic to the method used to construct them.
@@ 4 Experiments
The success of the inference methods was evaluated by considering samples drawn from the posterior distributions generated by Split-BOLFI and BOLFI, and the samples returned by the rejection ABC algorithm. We considered the RMSEs (Root Mean Square Errors, averaged across random seeds i.e. the RMSE of each posterior sample from the corresponding true generative value), AMEs (Absolute Mean Error, averaged across random seeds i.e. the absolute deviation of the sample mean from the true generative value) of the samples evaluated against the true generative values, and the standard deviations (SDs, averaged across posterior runs, i.e. the standard deviation of the posterior samples, representing the breadth of the posterior distribution, not considering the true generative value) of the returned samples, all averaged across the 50 random seeds used, and across the independent subsets across dimensions. Definitions for each of these are provided in Table 11. Further, for the Split-BOLFI method we recorded the AMAPEs (Absolute Maximum a Posterior Error, averaged across random seeds i.e. the absolute deviation of the MAP estimate from the true generative value), indicating the accuracy of the maximum a posteriori estimate of the updated belief distribution, which should be independent of the tempering level. We do not compare directly with the exact posterior of an untempered likelihood-based Bayesian update: in the presence of misspecification, then the standard Bayesian update is no longer the ideal, and we do not consider recreating it exactly to be the ultimate inferential target.
@@ 4.1 Overdispersed Gaussian
The inference result statistics of the $\mu$ and $\sigma$ parameters for maximum number of acquisitions used, with observed data sets of 500 data points. are presented in Tables 12, 13 and 14. The inference result statistics of the $\mu$ and $\sigma$ parameters for maximum number of acquisitions used, with observed data sets of 5000 data points, are presented in Tables 1, 2 and 3. The sizes of these datasets were chosen to demonstrate the behavior of the inference with the summary statistics exhibiting varying levels of stability, and do not necessarily reflect real-world investigative conditions.
@@
In Tables 12, 13 and 14, we see that in the well-specified case the point estimate provided by SB performs comparably well with BOLFI and the ABC estimate performed used much more simulations. The BOLFI algorithm loses accuracy substantially in the 10-dimensional parameter space compared with ABC and SB. We see that the SB method assigns more breadth to the posterior than the other methods when successful, especially when the kurtosis statistics is included, resulting in larger SD and RMSE measures. This is not necessarily appropriate, as the model is well-specified in this case, but may simply indicate that $n_{obs}=500$ is insufficiently large to reject the possibility of misspecification, and consequently some uncertainty is added. Comparing with the misspecified case, the results are substantially similar to the well-specified case except that SB with kurtosis adds substantially more posterior breadth having clearly identified the misspecification present. None of the other results change substantially, which is not necessarily desirable if we would like to identify misspecification and add posterior uncertainty when it is present. We see corresponding results as a function of the number of simulators acquisitions in Figure 9.
@@ 4.2.1 Simulator Study
The results for the simulator study are presented in Table 4 and Figure 6. We see the slightly counterintuitive result that some of the AME for BOLFI and Split-BOLFI are smaller in the misspecified case, but this is because the misspecified model is a lower-dimensional inference task. We observe that the AMEs for all of the parameters in the well-specified case are smaller for Split-BOLFI than for BOLFI, suggesting that the modularised approach has helped negotiate the higher-dimensional space. We do not observe a larger posterior breadth for Split-BOLFI when misspecification is introduced, but since the misspecification is brought about by a substantial change in the statistical model rather than the true generative process, then these posterior statistics are not necessarily comparable. Figure 6 suggests that Split-BOLFI makes constructive use of increasing number of acquisitions. BOLFI appears to become unstable with increasing number of simulations in the higher-dimensional case, which is not evident for the Split-BOLFI algorithm.
@@ Acknowledgements
Table 11: Definitions of the various performance metrics used. Performance metrics were computed for each random seed, then means and sds were computed across random seeds. AMAPE was only computed for Split-BOLFI, as this was the only method for which $\theta_{MAP}^{i}$ were recorded. Posterior samples $\theta_{j}^{i}$ are indexed by $n_{i}$ random seeds $i$ and $n_{j}$ within-posterior sample indices $j$. True generative values $\theta_{T}^{i}$ are indexed by random seed $i$.

### v5

@@ Abstract
Likelihood-free inference for simulator-based statistical models has developed rapidly from its infancy to a useful tool for practitioners. However, models with more than a handful of parameters still generally remain a challenge for the Approximate Bayesian Computation (ABC) based inference. To advance the possibilities for performing likelihood-free inference in higher dimensional parameter spaces, we introduce an extension of the popular Bayesian optimisation based approach to approximate discrepancy functions in a probabilistic manner which lends itself to an efficient exploration of the parameter space. Our approach achieves computational scalability for higher dimensional parameter spaces by using separate acquisition functions, discrepancies, and associated summary statistics for distinct subsets of the parameters. The efficient additive acquisition structure is combined with exponentiated loss-likelihood to provide a misspecification-robust characterisation of posterior distributions for subsets of model parameters. The method successfully performs computationally efficient inference in a moderately sized parameter space and compares favourably to existing modularised ABC methods. We further illustrate the potential of this approach by fitting a bacterial transmission dynamics model to a real data set, which provides biologically coherent results on strain competition in a 30-dimensional parameter space.
@@ 3.2 Bayesian Optimization
Each GP model trained for each discrepancy component has an analytically accessible predictive mean and standard deviation $\mu_{j}^{*}([\theta^{*}]_{j})$ and $\sigma_{j}^{*}([\theta^{*}]_{j})$, calculated from the model components and observed data . Through the use of the Lower Confidence Bound (LCB) acquisition function $a_{LCB}()$ (Srinivas et al., 2010) with an exploration-exploitation tradeoff parameter $\beta$, this gives rise to the following acquisition function minimum for the $j$th subset:
@@
Other acquisition functions can be used in addition to the LCB, including Expected Improvement (EI) (Jones et al., 1998), Probability of Improvement (PI) (Kushner, 1964), and those specifically designed for likelihood-free inference (Järvenpää et al., 2019a). LCB comes with convergence guarantees under some conditions (Srinivas et al., 2010). The parameter $\beta$ in the LCB determines the compromise between exploration and exploitation in the optimization routine.Other acquisitions functions, including EI and PI, use similar hyperparameters for negotiating the tradeoff of exploration with exploitation (Gan et al., 2021). For researchers looking to derive an appropriate $\beta$ parameter value for their application can follow the principle that larger parameter values emphasizing exploration do not take advantage of the current surrogate fit but evaluate more evenly in the whole optimization space, and smaller parameter values emphasizing exploitation focus the acquisition locations where surrogate fit is already optimal (Brochu et al., 2010).
In the analyses in this article, we performed preliminary inference runs with varying values of $\beta$, intended to determine the appropriate value of $\beta$ for the full inference procedure. We initially tried values of several negative powers of 10, i.e. [0.001, 0.01, 0.1, 1.], and examined the posterior distributions produced to determine whether the results of inference were satisfactory. These preliminary runs were performed with simulated data drawn with known generative parameter values, allowing comparison with the underlying true values. Overly small values of $\beta$ (0.01 or smaller) could result in either convergence to a spurious local posterior mode, or overly aggressive concentration to the true posterior mode, with inadequate characterization of the rest of the posterior. Conversely, overly large values of $\beta$ resulted in a much more even selection of acquisitions within the parameter bounds, at the potential expense of thoroughly investigating, or sometimes even identifying, the true posterior mode. Given the low dimensionality of the individual subsets used in Split-BOLFI, such phenomena could be identified through diagrams such as those presented in Figure 4, although quantitative metrics could also be used to evaluate the success of inference. Based on our exploratory analyses of the problems considered here we have fixed parameter $\beta=0.1$. This choice may not necessarily generalize to other models and data sets, and a wider exploration is recommended for each new application.
For users of Split-BOLFI looking to apply the methodology to new problems, we advocate a similar approach of pursuing preliminary runs with varying values of $\beta$ to establish the resulting posterior characteristics. The value of $\beta=0.1$ that we used in this article is by no means a universal solution for every simulator, but may represent a reasonable place to start an exploration of $\beta$ for preliminary analysis. The optimal value of $\beta$ will depend on the posterior geometry (conditional on choice of summary statistics), computational budget, and inferential goals. The final decision of selecting $\beta$ rests with the practising statistician and their priorities: for a given acquisition budget, smaller values of $\beta$ will result in a more thorough characterization of the posterior mode compared to the rest of the distribution (with the risk of getting caught in a local mode), whereas larger values of $\beta$ will result in a more thorough characterization of the entire posterior compared to the mode (with the risk of missing the global posterior mode entirely). The former may be preferred when a posterior mode with some degree of uncertainty is desired, whereas the latter may be preferred when full uncertainty propagation from the parameter space is the goal.
@@
i.e. whichever is larger of the minimum GP predictive mean value $\mu_{j}^{*}([\theta^{*}]_{j})$ and the minimum of the observed discrepancy values $d^{obs}_{j}$ from the simulator. Because the GP function is defined over all real numbers, the predictive mean can sometimes dip below zero, making $\underset{[\theta^{*}]_{j}}{\min}(\mu_{j}^{*}([\theta^{*}]_{j}))$ hard to interpret, so in these situations, we use the minimum observed discrepancy value $\min(d^{obs}_{j})$, as this is the smallest empirically validated quantity that will lead to numerically stable solutions. It is also possible to use a non-linear transformation (e.g. logarithmic) to avoid this problem: as explored in (Gutmann and Corander, 2016) it can help in some instances of numerical instability and misspecification of the discrepancy model, although it was less influential than the introduction of noise into the acquisition process.
@@ 3.4 Summary Statistics and Partitioning
We encourage the use of any of the above methods, or in some combination, to help derive sets of summary statistics appropriate for inference. Establishing a successful Bayesian workflow generally relies upon an iterative process of predictive checks, computation validation, and model evaluation, among other steps (Gelman et al., 2020), with “black box” inference solutions often falling outside of this paradigm. (Brynjarsdóttir and O’Hagan, 2014) recommend further: “We believe that formulating prior information about model discrepancy is the key to realistic calibration. In most cases, modelers have some idea about what physics or processes the simulator is missing, and we need to use the best judgments of the modelers and model users about how the simulator’s deficiencies will translate into model discrepancy.” For the examples in this article, we have used summaries derived from our understanding of the simulators, but the inference methodology is agnostic to the method used to construct them.
@@ 4 Experiments
The success of the inference methods was evaluated by considering samples drawn from the posterior distributions generated by Split-BOLFI and BOLFI, and the samples returned by the rejection ABC algorithm. We considered the RMSEs (Root Mean Square Errors, averaged across random seeds i.e. the RMSE of each posterior sample from the corresponding true generative value), AMEs (Absolute Mean Error, averaged across random seeds i.e. the absolute deviation of the sample mean from the true generative value) of the samples evaluated against the true generative values, standard deviations (SDs, averaged across posterior runs, i.e. the standard deviation of the posterior samples, representing the breadth of the posterior distribution, not considering the true generative value) of the returned samples, and 50% posterior coverage of the true generative value compared with the first and third posterior quartiles, all averaged across the 50 random seeds used, and across the independent subsets across dimensions. Definitions for each of these are provided in Table 12. Further, for the Split-BOLFI method we recorded the AMAPEs (Absolute Maximum a Posterior Error, averaged across random seeds i.e. the absolute deviation of the MAP estimate from the true generative value), indicating the accuracy of the maximum a posteriori estimate of the updated belief distribution, which should be independent of the tempering level. We do not compare directly with the exact posterior of an untempered likelihood-based Bayesian update: in the presence of misspecification, then the standard Bayesian update is no longer the ideal, and we do not consider recreating it exactly to be the ultimate inferential target.
@@ 4.1 Overdispersed Gaussian
The inference result statistics of the $\mu$ and $\sigma$ parameters for maximum number of acquisitions used, with observed data sets of 500 data points. are presented in Tables 13, 14, 15 and 16. The inference result statistics of the $\mu$ and $\sigma$ parameters for maximum number of acquisitions used, with observed data sets of 5000 data points, are presented in Tables 1, 2, 3 and 17. The sizes of these datasets were chosen to demonstrate the behavior of the inference with the summary statistics exhibiting varying levels of stability, and do not necessarily reflect real-world investigative conditions.
@@
In Tables 13, 14 and 15, we see that in the well-specified case the point estimate provided by SB performs comparably well with BOLFI and the ABC estimate performed used much more simulations. The BOLFI algorithm loses accuracy substantially in the 10-dimensional parameter space compared with ABC and SB. We see that the SB method assigns more breadth to the posterior than the other methods when successful, especially when the kurtosis statistics is included, resulting in larger SD and RMSE measures. This is not necessarily appropriate, as the model is well-specified in this case, but may simply indicate that $n_{obs}=500$ is insufficiently large to reject the possibility of misspecification, and consequently some uncertainty is added. Comparing with the misspecified case, the results are substantially similar to the well-specified case except that SB with kurtosis adds substantially more posterior breadth having clearly identified the misspecification present. None of the other results change substantially, which is not necessarily desirable if we would like to identify misspecification and add posterior uncertainty when it is present. We see corresponding results as a function of the number of simulators acquisitions in Figure 9.
We also investigated coverages and they indicated that the ABC methods generally report less than 50% coverage, whereas Split-BOLFI and BOLFI show generally larger than 50% coverage. We observe that the ABC methods contribute slightly larger coverage when the kurtosis statistic is introduced in the misspecified case, whereas the other methods’ coverages do not vary substantially with changes in experimental setup. Details can be found in Appendix Tables 16 and 17.
@@ 4.2.1 Simulator Study
Results for the simulator study are presented in Table 4 and Figure 6. We see the slightly counterintuitive result that some of the AME for BOLFI and Split-BOLFI are smaller in the misspecified case, but this is because the misspecified model is a lower-dimensional inference task. We observe that the AMEs for all of the parameters in the well-specified case are smaller for Split-BOLFI than for BOLFI, suggesting that the modularised approach has helped negotiate the higher-dimensional space. We do not observe a larger posterior breadth for Split-BOLFI when misspecification is introduced, but since the misspecification is brought about by a substantial change in the statistical model rather than the true generative process, then these posterior statistics are not necessarily comparable. Figure 6 suggests that Split-BOLFI makes constructive use of increasing number of acquisitions. BOLFI appears to become unstable with increasing number of simulations in the higher-dimensional case, which is not evident for the Split-BOLFI algorithm. The coverage analysis shows consistently greater than 50% coverage for Split-BOLFI, and conflicting results for BOLFI in the well and misspecified cases: details can be found in Appendix Table 18.
@@ Acknowledgements
ABC SB BOLFI
$p=2$ $n_{b}=1$ 0.24 (0.43) 0.82 (0.48) 0.64 (0.48)
$p=6$ $n_{b}=1$ 0.15 (0.35) 0.85 (0.35) 0.15 (0.35)
$n_{b}=2$ 0.14 (0.35) 0.69 (0.48) 0.35 (0.48)
$n_{b}=3$ 0.12 (0.33) 0.83 (0.48) 0.37 (0.48)
$p=2$ $n_{b}=1$ 0.25 (0.44) 0.83 (0.46) 0.7 (0.46)
$p=6$ $n_{b}=1$ 0.19 (0.39) 0.72 (0.4) 0.19 (0.4)
$n_{b}=2$ 0.18 (0.38) 0.84 (0.48) 0.37 (0.48)
$n_{b}=3$ 0.19 (0.4) 0.85 (0.49) 0.42 (0.49)
Table 11: Results for a pairwise blocky Cholesky precision model, in the well-specified case (above double line) and misspecified case (below double line). We consider data dimensionalities $p$ equal to 2 and 6, producing 1 or 3 blocks in each case, indexed by $n_{b}$. Algorithms are ABC, Split-BOLFI, and BOLFI. Results of the mean(sd) of 50% coverage with a data set of 500 observations. Error is based on the point estimate which was sample mean for ABC and BOLFI and MAP for Split-BOLFI.
@@
The 50% coverage metric in Table 11 suggests that ABC shows consistently lower than 50% coverage, Split-BOFLI exhibits somewhat higher than 50% coverage, and BOLFI exhibits lower than 50% coverage on the high dimensional problem, and higher than 50% coverage on the lower-dimensional problem. We do not observe substantial variation in coverage in any of the methods between the well and misspecified conditions.
@@
Table 12: Definitions of the various performance metrics used. Performance metrics were computed for each random seed, then means and sds were computed across random seeds. AMAPE was only computed for Split-BOLFI, as this was the only method for which $\theta_{MAP}^{i}$ were recorded. Posterior samples $\theta_{j}^{i}$ are indexed by $n_{i}$ random seeds $i$ and $n_{j}$ within-posterior sample indices $j$. True generative values $\theta_{T}^{i}$ are indexed by random seed $i$. $Q1[\theta_{j}^{i}]$ and $Q3[\theta_{j}^{i}]$ are respectively the first and third quartiles of the $i$th posterior sample $\theta_{j}^{i}$.
@@
$\mu_{1}$ 0 (0.05) 0 (0) 0.94 (0.24) 0.98 (0.14) 0.74 (0.44)
$\sigma_{1}$ 0.01 (0.08) 0.01 (0.07) 1 (0) 1 (0) 0.92 (0.27)
$\mu_{5}$ 0 (0.05) 0 (0.05) 0.94 (0.25) 0.99 (0.09) 0.5 (0.5)
$\sigma_{5}$ 0.01 (0.07) 0.01 (0.07) 0.99 (0.09) 1 (0) 0.64 (0.48)
$\mu_{1}$ 0 (0.03) 0 (0.05) 0.94 (0.24) 0.96 (0.2) 0.68 (0.47)
$\sigma_{1}$ 0 (0.05) 0.01 (0.1) 0.94 (0.24) 0.98 (0.14) 0.76 (0.43)
$\mu_{5}$ 0 (0.05) 0 (0.04) 0.94 (0.23) 0.95 (0.22) 0.49 (0.5)
$\sigma_{5}$ 0.01 (0.08) 0.01 (0.08) 0.89 (0.31) 1 (0.06) 0.65 (0.48)
Table 16: Results for a Gaussian statistical model with parameters $\mu_{p}$ and $\sigma_{p}$ in the well-specified case (above line) and misspecified case (below line) for $p=1$ and $p=5$. Algorithms are ABC, Split-BOLFI, and BOLFI using mean and standard deviation summaries $m,s$ and additionally kurtosis summary $m,s,k$. Results of the mean(sd) of 50% coverage with a data set of 500 observations. Error is based on the point estimate which was sample mean for ABC and BOLFI and MAP for Split-BOLFI.
$\mu_{1}$ 0 (0) 0 (0.05) 1 (0) 1 (0) 0.92 (0.27)
$\sigma_{1}$ 0.01 (0.09) 0.01 (0.07) 1 (0) 1 (0) 1 (0)
$\mu_{5}$ 0 (0.02) 0 (0.03) 1 (0.06) 1 (0) 0.53 (0.5)
$\sigma_{5}$ 0 (0.06) 0.01 (0.07) 1 (0) 1 (0) 0.63 (0.48)
$\mu_{1}$ 0 (0.06) 0.05 (0.22) 1 (0) 0.96 (0.2) 0.96 (0.2)
$\sigma_{1}$ 0.01 (0.07) 0.05 (0.22) 1 (0) 1 (0) 0.9 (0.3)
$\mu_{5}$ 0 (0.04) 0.05 (0.22) 0.99 (0.09) 0.94 (0.23) 0.6 (0.49)
$\sigma_{5}$ 0.01 (0.08) 0.05 (0.22) 1 (0) 0.99 (0.09) 0.57 (0.5)
Table 17: Results for a Gaussian statistical model with parameters $\mu_{p}$ and $\sigma_{p}$ in the well-specified case (above line) and misspecified case (below line) for $p=1$ and $p=5$. Algorithms are ABC, Split-BOLFI, and BOLFI using mean and standard deviation summaries $m,s$ and additionally kurtosis summary $m,s,k$. Results of the mean(sd) of 50% coverage with a data set of 5000 observations. Error is based on the point estimate which was sample mean for ABC and BOLFI and MAP for Split-BOLFI.
BOLFI SB
$\beta$ 0 (0) 1 (0)
$\Lambda$ 0.44 (0.5) 1 (0)
Avg. $\theta$ 0 (0) 0 (0)
$\beta$ 0.8 (0.4) 0.96 (0.2)
$\Lambda$ 0.74 (0.44) 1 (0)
Avg. $\theta$ 0 (0) 0 (0)
Table 18: Results for a daycare simulator model in the well-specified (above double line) and misspecified case (below double line) i.e with a high-dimensional and low-dimensional statistical model and generative models, respectively. Results are presented as mean(sd) of the 50% coverage for BOLFI and Split-BOLFI for the parameters $\beta$, $\Lambda$ and averaged over $\theta$. The coverage for the $\theta$ parameter is not meaningful since it lies on the boundary of the parameter space